bartja83:

Find the magnitude of the vector A⃗ = (5.5 m ) xhat + (-2.2 m ) yhat.

snowie:

@oliver69

toga:

To find the magnitude of a vector, we use the formula: |A⃗| = sqrt(Ax^2 + Ay^2) Where Ax and Ay are the x and y components of the vector A⃗. For the given vector A⃗ = (5.5 m ) xhat + (-2.2 m ) yhat, Ax = 5.5 m and Ay = -2.2 m. So, the magnitude of the vector A⃗ is: |A⃗| = sqrt(5.5^2 + (-2.2)^2) m |A⃗| = sqrt(30.25 + 4.84) m |A⃗| = sqrt(35.09) m |A⃗| = 5.92 m (approx) Hence, the magnitude of the vector A⃗ is 5.92 meters (approx).

arissperf:

this math is too hard just say ur dog ate ur computer duhhhh